THE PERIMETER OF A RECTANGLE IS 34 CM. HOW DO YOU SOLVE FOR THE LENGTH AND HEIGHT?
Answer by richard1234(7193)

We have 2(L+H) = 34 --> L+H = 17. So L = 17-H, H = 17-L.